# Unilaw-R1: A Large Language Model for Legal Reasoning with Reinforcement Learning and Iterative Inference

### v1

@@ Abstract
Reasoning-focused large language models (LLMs) are rapidly evolving across various domains, yet their capabilities in handling complex legal problems remains underexplored. In this paper, we introduce Unilaw-R1, a large language model tailored for legal reasoning. With a lightweight 7-billion parameter scale, Unilaw-R1 significantly reduces deployment cost while effectively tackling three core challenges in the legal domain: insufficient legal knowledge, unreliable reasoning logic, and weak business generalization. To address these issues, we first construct Unilaw-R1-Data, a high-quality dataset containing $\sim$17K distilled and screened chain-of-thought (CoT) samples. Based on this, we adopt a two-stage training strategy combining Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), which significantly boosts the model’s performance on complex legal reasoning tasks and supports interpretable decision-making in legal AI applications. To assess legal reasoning ability, we also introduce Unilaw-R1-Eval, a dedicated benchmark designed to evaluate models across single- and multi-choice legal tasks. Unilaw-R1 demonstrates strong results on authoritative benchmarks, outperforming all models of similar scale and achieving performance on par with the much larger DeepSeek-R1-Distill-Qwen-32B (54.9%). Following domain-specific training, it also showed significant gains on LawBench and LexEval, exceeding Qwen-2.5-7B-Instruct (46.6%) by an average margin of 6.6%. Code is available at: https://github.com/Hanscal/Unilaw-R1. Corresponding authors: Hua Cai (hua.cai@unidt.com)
and Xuli Shen (xlshen20@fudan.edu.cn)
Accepted by EMNLP 2025.
@@
Hua Cai1 $\dagger$, Shuang Zhao1, Liang Zhang1, Xuli Shen1,2 $\dagger$, Qing Xu1, Weilin Shen1, Zihao Wen2, Tianke Ban2 1UniDT, 2Fudan University
@@ 1 Introduction
In recent years, the rapid iteration of large language models (LLMs) has significantly propelled the evolution of artificial intelligence towards artificial general intelligence (AGI). Models such as OpenAI’s o1-series (OpenAI Team, 2024) have enhanced their ability for complex reasoning tasks by extending the length of the "chain-of-thought" through an "exploration-reflection-iteration" mechanism. Similar o1-like LLMs, such as QwQ (Qwen, 2025) and Marco-o1 (Zhao et al., 2024b), have demonstrated significant improvements across tasks like mathematics, programming, and logical reasoning. Although general reasoning models exhibit considerable potential, their application in specialized domains such as legal is limited. Legal reasoning requires not only legal, economic and mathematical knowledge, but also step-by-step and verifiable logic. Existing models face three major challenges: (1) inconsistencies in legal data increase preprocessing complexity and weaken reasoning (Koenecke et al., 2025; Mishra et al., 2025; Sheik et al., 2024; Steging et al., 2023; Aumiller et al., 2021); (2) the black-box nature of LLMs lack transparency, falling short of traceability standards (Wang et al., 2023; Zhao et al., 2024a; Tong et al., 2024; Chaudhary, 2024); and (3) insufficient legal knowledge leads to unreliable or incoherent reasoning processes (Blair-Stanek and Van Durme, 2025; Dahl et al., 2024). Moreover, effective legal reasoning must adhere to both the external validity of codified law and the internal procedural consistency that ensures fairness and predictability in legal interpretation (Zou, 2021; Raz, 2009; Fuller, 1969).
@@ 2 Related Work
The capabilities of large language models have advanced rapidly through innovations in training paradigms and reasoning strategies. The o1-series models (Jaech et al., 2024) introduced iterative "exploration-reflection" mechanisms that lengthen the CoT process, thereby improving reasoning depth. Subsequent efforts such as QwQ (Qwen, 2025), Marco-o1 (Zhao et al., 2024b) and Fin-R1 (Liu et al., 2025) extended this approach across domains including logic, mathematics, and finance. In the legal domain, compliant adaptations of o1-class models, such as HK-O1aw and PatientSeek (HKAIR, 2024; whyhow ai, 2025), have shown the potential of LLMs in simulating human-like legal reasoning. Yu et al. Yu et al. (2025) further pushed this frontier by employing test-time scaling techniques to enhance performance on legal tasks.
Distinct from the above, DeepSeek-R1 (Guo et al., 2025) takes an efficient reinforcement learning (RL) approach, training LLMs via thousands of steps of unsupervised RL combined with a cold-start corpus and multi-stage curriculum learning. This strategy results in emergent reasoning capabilities and improved readability, highlighting the promise of RL in scaling inference power.
@@ 3.1 Overview
We propose a two-stage framework for legal reasoning model construction, as illustrated in Figure 1. In the data generation stage, we construct a high-quality legal reasoning dataset, Unilaw-R1-Data, by leveraging a data distillation approach grounded in DeepSeek-R1 and incorporating an LLM-as-judge filtering mechanism (Xu et al., 2023) to ensure annotation consistency and reasoning rigor. In the model training stage, we develop the Unilaw-R1 model based on Qwen2.5-7B-Instruct (Yang et al., 2024), utilizing Supervised Fine-Tuning (SFT) in combination with the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024). To further enhance reasoning performance, we introduce an iterative inference mechanism with a collaborative Assessor-Reviser agent setup, enabling the model to refine its reasoning trajectory for more accurate, coherent, and legally sound outputs. The overall process ensures that the model delivers structured, standardized outputs aligned with professional requirements.
@@ 3.2.1 Data Source
Unilaw-R1-Data consists of objective question answering entries in the legal domain, drawn from two primary sources: the open-source JEC-QA dataset (Yue et al., 2023) and proprietary data. JEC-QA includes 26, 365 multiple-choice questions, each with a question and four options. The proprietary portion includes 1, 700 multiple-choice question answering entries from National Judicial Examination of China from year 2015 to 2021. These were collected as PDFs, converted to markdown using Mineru (Wang et al., 2024), and structured into question-answer pairs via regex-based extraction. All entries were manually reviewed for accuracy. From our proprietary data, 800 high-quality samples were retained to form the Unilaw-R1-Eval set for model evaluation.
@@ 3.2.2 Data Processing
All generated chains from Chain Rewriting and Explanation Generation modules, along with those filtered by Answer Check module, are passed into the Reasoning Selection phase to evaluate the plausibility and legal soundness of multiple reasoning trajectories using the instruction model DeepSeek-V3 (Liu et al., 2024). Responses are scored based on their adherence to legal reasoning principles, such as the correct application of rules, consistency with precedent, and logical coherence. These dimensions were employed to comprehensively evaluate the model’s reasoning trajectory data. The model’s reasoning path must not only lead to the correct answer but also demonstrate a valid and interpretable argumentative structure. When multiple valid paths exist, we prioritize those that align more closely with recognized legal standards and practices. Further details on the experimental setup and findings are provided in Appendix A.
@@ 4.1 Datasets
We evaluate our model on Unilaw-R1-Eval dataset and two additional Chinese legal domain multi-task benchmarks: LawBench (Fei et al., 2024) dataset and LexEval (Li et al., 2024). LawBench assesses the legal capabilities of LLMs across three cognitive levels: memory, understanding, and application. It comprises 20 tasks with various formats, including multiple-choice, extraction, generation, and regression, simulating real-world legal scenarios such as statute prediction, case analysis, and legal consultation. LexEval, the largest and most comprehensive Chinese legal benchmarking dataset, evaluates performance of LLMs across six cognitive abilities defined by the LexCog taxonomy: memory, understanding, logical reasoning, discrimination, generation, and ethics. It consists of 14,150 entries across 23 legal tasks, providing a diverse set for evaluating LLM performance.
@@ 4.4 Ablation Study
As shown in Table 3, we compared zero-shot CoT (Wei et al., 2022), best-of-$k$ sampling, majority vote (Wang et al., ) and iterative inference methods across single-choice (SC) and multi-choice (MC) tasks. The zero-shot CoT baseline achieved 53.8% accuracy on SC tasks and 23.2% on MC tasks. Implementing best-of-$k$ ($k=10$) sampling and majority vote led to improvements, raising the average accuracy from 39.5% to 45.2% and 45.7%, respectively. The iterative inference approach demonstrated more substantial gains. With a single iteration, SC accuracy increased to 65.5% and MC to 33.4%. The performance gains form further iterations were limited: the second iteration achieved 66.3% accuracy on SC and 33.8% on MC, while the third iteration reached 65.7% (SC) and 34.3% (MC), respectively.

### v2

@@ Abstract
Reasoning-focused large language models (LLMs) are rapidly evolving across various domains, yet their capabilities in handling complex legal problems remains underexplored. In this paper, we introduce Unilaw-R1, a large language model tailored for legal reasoning. With a lightweight 7-billion parameter scale, Unilaw-R1 significantly reduces deployment cost while effectively tackling three core challenges in the legal domain: insufficient legal knowledge, unreliable reasoning logic, and weak business generalization. To address these issues, we first construct Unilaw-R1-Data, a high-quality dataset containing $\sim$17K distilled and screened chain-of-thought (CoT) samples. Based on this, we adopt a two-stage training strategy combining Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), which significantly boosts the model’s performance on complex legal reasoning tasks and supports interpretable decision-making in legal AI applications. To assess legal reasoning ability, we also introduce Unilaw-R1-Eval, a dedicated benchmark designed to evaluate models across single- and multi-choice legal tasks. Unilaw-R1 demonstrates strong results on authoritative benchmarks, outperforming all models of similar scale and achieving performance on par with the much larger DeepSeek-R1-Distill-Qwen-32B (54.9%). Following domain-specific training, it also showed significant gains on LawBench and LexEval, exceeding Qwen-2.5-7B-Instruct (46.6%) by an average margin of 6.6%. Code is available at: https://github.com/Hanscal/Unilaw-R1. Corresponding authors: Tianke Ban (bantianke@fudan.
edu.cn), Hua Cai (hua.cai@unidt.com), Xuli Shen (xlshen20@fudan.edu.cn).
@@
Hua Cai1 $\dagger$, Shuang Zhao1, Liang Zhang1, Xuli Shen1,2 $\dagger$, Qing Xu1, Weilin Shen1, Zihao Wen2, Tianke Ban2 $\dagger$ 1UniDT, 2Fudan University
@@ 1 Introduction
In recent years, the rapid iteration of large language models (LLMs) has significantly propelled the evolution of artificial intelligence towards artificial general intelligence (AGI). Models such as OpenAI’s o1-series (openai2024llms) have enhanced their ability for complex reasoning tasks by extending the length of the "chain-of-thought" through an "exploration-reflection-iteration" mechanism. Similar o1-like LLMs, such as QwQ (qwen) and Marco-o1 (zhao2024marco), have demonstrated significant improvements across tasks like mathematics, programming, and logical reasoning. Although general reasoning models exhibit considerable potential, their application in specialized domains such as legal is limited. Legal reasoning requires not only legal, economic and mathematical knowledge, but also step-by-step and verifiable logic. Existing models face three major challenges: (1) inconsistencies in legal data increase preprocessing complexity and weaken reasoning (koenecke2025tasks; mishra2025investigating; sheik2024neural; steging2023improving; aumiller2021structural); (2) the black-box nature of LLMs lack transparency, falling short of traceability standards (wang2023alpha; zhao2024explainability; tong2024ploutos; chaudhary2024unveiling); and (3) insufficient legal knowledge leads to unreliable or incoherent reasoning processes (blair2025llms; dahl2024large). Moreover, effective legal reasoning must adhere to both the external validity of codified law and the internal procedural consistency that ensures fairness and predictability in legal interpretation (zou2021nine; raz2009authority; fuller1969morality).
@@ 2 Related Work
The capabilities of large language models have advanced rapidly through innovations in training paradigms and reasoning strategies. The o1-series models (jaech2024openai) introduced iterative "exploration-reflection" mechanisms that lengthen the CoT process, thereby improving reasoning depth. Subsequent efforts such as QwQ (qwen), Marco-o1 (zhao2024marco) and Fin-R1 (liu2025fin) extended this approach across domains including logic, mathematics, and finance. In the legal domain, compliant adaptations of o1-class models, such as HK-O1aw and PatientSeek (HK-O1aw; PatientSeek), have shown the potential of LLMs in simulating human-like legal reasoning. Yu et al. yu2025evaluating further pushed this frontier by employing test-time scaling techniques to enhance performance on legal tasks.
Distinct from the above, DeepSeek-R1 (guo2025deepseek) takes an efficient reinforcement learning (RL) approach, training LLMs via thousands of steps of unsupervised RL combined with a cold-start corpus and multi-stage curriculum learning. This strategy results in emergent reasoning capabilities and improved readability, highlighting the promise of RL in scaling inference power.
@@ 3.1 Overview
We propose a two-stage framework for legal reasoning model construction, as illustrated in Figure 1. In the data generation stage, we construct a high-quality legal reasoning dataset, Unilaw-R1-Data, by leveraging a data distillation approach grounded in DeepSeek-R1 and incorporating an LLM-as-judge filtering mechanism (xu2023mmbenchmark) to ensure annotation consistency and reasoning rigor. In the model training stage, we develop the Unilaw-R1 model based on Qwen2.5-7B-Instruct (yang2024qwen2), utilizing Supervised Fine-Tuning (SFT) in combination with the Group Relative Policy Optimization (GRPO) algorithm (shao2024deepseekmath). To further enhance reasoning performance, we introduce an iterative inference mechanism with a collaborative Assessor-Reviser agent setup, enabling the model to refine its reasoning trajectory for more accurate, coherent, and legally sound outputs. The overall process ensures that the model delivers structured, standardized outputs aligned with professional requirements.
@@ 3.2.1 Data Source
Unilaw-R1-Data consists of objective question answering entries in the legal domain, drawn from two primary sources: the open-source JEC-QA dataset (yue2023disc) and proprietary data. JEC-QA includes 26, 365 multiple-choice questions, each with a question and four options. The proprietary portion includes 1, 700 multiple-choice question answering entries from National Judicial Examination of China from year 2015 to 2021. These were collected as PDFs, converted to markdown using Mineru (wang2024mineruopensourcesolutionprecise), and structured into question-answer pairs via regex-based extraction. All entries were manually reviewed for accuracy. From our proprietary data, 800 high-quality samples were retained to form the Unilaw-R1-Eval set for model evaluation.
@@ 3.2.2 Data Processing
All generated chains from Chain Rewriting and Explanation Generation modules, along with those filtered by Answer Check module, are passed into the Reasoning Selection phase to evaluate the plausibility and legal soundness of multiple reasoning trajectories using the instruction model DeepSeek-V3 (liu2024deepseek). Responses are scored based on their adherence to legal reasoning principles, such as the correct application of rules, consistency with precedent, and logical coherence. These dimensions were employed to comprehensively evaluate the model’s reasoning trajectory data. The model’s reasoning path must not only lead to the correct answer but also demonstrate a valid and interpretable argumentative structure. When multiple valid paths exist, we prioritize those that align more closely with recognized legal standards and practices. Further details on the experimental setup and findings are provided in Appendix A.
@@ 4.1 Datasets
We evaluate our model on Unilaw-R1-Eval dataset and two additional Chinese legal domain multi-task benchmarks: LawBench (fei2024lawbench) dataset and LexEval (li2024lexeval). LawBench assesses the legal capabilities of LLMs across three cognitive levels: memory, understanding, and application. It comprises 20 tasks with various formats, including multiple-choice, extraction, generation, and regression, simulating real-world legal scenarios such as statute prediction, case analysis, and legal consultation. LexEval, the largest and most comprehensive Chinese legal benchmarking dataset, evaluates performance of LLMs across six cognitive abilities defined by the LexCog taxonomy: memory, understanding, logical reasoning, discrimination, generation, and ethics. It consists of 14,150 entries across 23 legal tasks, providing a diverse set for evaluating LLM performance.
@@ 4.4 Ablation Study
As shown in Table 3, we compared zero-shot CoT (wei2022chain), best-of-$k$ sampling, majority vote (wangself) and iterative inference methods across single-choice (SC) and multi-choice (MC) tasks. The zero-shot CoT baseline achieved 53.8% accuracy on SC tasks and 23.2% on MC tasks. Implementing best-of-$k$ ($k=10$) sampling and majority vote led to improvements, raising the average accuracy from 39.5% to 45.2% and 45.7%, respectively. The iterative inference approach demonstrated more substantial gains. With a single iteration, SC accuracy increased to 65.5% and MC to 33.4%. The performance gains form further iterations were limited: the second iteration achieved 66.3% accuracy on SC and 33.8% on MC, while the third iteration reached 65.7% (SC) and 34.3% (MC), respectively.